OpenStudy (anonymous):

If I am given two functions with square roots in them, am I allowed to square the function before I apply l'Hopital's Rule or not?

OpenStudy (jhannybean):

Give an exampe

OpenStudy (anonymous):

For example, in this problem (I already put it in limit form) lim(x->infinity) sqrtx/sqrt(x+1)

OpenStudy (anonymous):

Am I allowed to square the whole thing before I l'Hopital?

OpenStudy (anonymous):

If I take the square root of my answer anyways

OpenStudy (vera_ewing):

Do you know how to start this?

OpenStudy (jhannybean):

\[\lim_{x\rightarrow \infty} \frac{\sqrt{x}}{\sqrt{x+1}} \] I would just use LH rule right away. without changing it.

OpenStudy (freckles):

no you can't just square the whole function

OpenStudy (jhannybean):

you could square the entire function if it was one function.... but here you are dealing with 2 functions.

OpenStudy (anonymous):

Why not? My teacher said that we could, but she didn't say anything about before LH

OpenStudy (anonymous):

As long as we do the opposite to our answer

OpenStudy (freckles):

\[\lim_{x \rightarrow 3}x=3 \\ \text{ but } \lim_{x \rightarrow 3}x^2=9 \\ 9 \neq 3\]

OpenStudy (anonymous):

I'm comparing rates of growth

OpenStudy (anonymous):

Freckles, I'm not setting the limit equal to anything

OpenStudy (anonymous):

I just have to prove that the limit comes out to a constant other than 0

OpenStudy (freckles):

\[\sqrt{ \lim_{x \rightarrow \infty} \frac{x}{x+1}}\] you don't need l'hospital

OpenStudy (freckles):

the inside function goes to what ?

OpenStudy (anonymous):

Well isn't that the same thing as if I squared the composed functions and took the square root of my answer?

OpenStudy (jhannybean):

\[\lim_{x\rightarrow \infty} \frac{x^{\color{red}{1}}}{x^{\color{red}{1}}+1}= ~? \]

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Then you take the square root and get 1

OpenStudy (freckles):

you can't say this: \[\lim_{x \rightarrow a}f(x) =\lim_{x \rightarrow a}(f(x))^2 \\ \text{ but you can say this } \\ \lim_{x \rightarrow a} f(x) =\lim_{x \rightarrow a} \sqrt{f(x))^2} \text{ provided } f(x)>0\]

OpenStudy (anonymous):

No, the square root of the LIMIT

OpenStudy (anonymous):

the whole thing

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}f(x)=\sqrt{\lim_{x \rightarrow \infty} f(x)^{2}}\]

OpenStudy (anonymous):

Why can't I do that

OpenStudy (freckles):

I didn't say you couldn't do that

OpenStudy (anonymous):

Thats what I meant

OpenStudy (jhannybean):

That's legit.

OpenStudy (freckles):

oh i thought you just meant square the whole thing and be done with it

OpenStudy (freckles):

lol

OpenStudy (anonymous):

I square the composed functions and take the square root of the answer xD

OpenStudy (jhannybean):

Same, freckles.

OpenStudy (anonymous):

Sorry!!!

OpenStudy (jhannybean):

Yes, and because \(x\rightarrow \infty\) it works.

OpenStudy (anonymous):

Ok! So I can square before I use l'hopital?

OpenStudy (freckles):

but I think the problem is easier if we just look at it like this: \[\lim_{x \rightarrow \infty}\sqrt{\frac{x}{x+1}}=\sqrt{\lim_{x \rightarrow \infty} \frac{x}{x+1}}\] and just find the limit inside

OpenStudy (freckles):

and yes you can use l'hopistal to find \[\lim_{x \rightarrow \infty}\frac{x}{x+1}\] but it isn't required

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}\frac{1}{1+\frac{1}{x}}=\frac{1}{1+0}=\frac{1}{1}=1\]

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}\sqrt{\frac{x}{x+1}}=\sqrt{\lim_{x \rightarrow \infty} \frac{x}{x+1}}=\sqrt{1}\]

OpenStudy (anonymous):

OH OK. Thank you guys so much!!

OpenStudy (jhannybean):

I didn't know you could take the limit inside a function like that.